you need to put floor covering in bedroom1 nd 2. how much floor covering in squre metre will you need?
length of room 1 is 5.0m and width is 5.0
lenth of room 2 is 4.0m and width is 5.0

Answer by nerdybill(7384)

you need to put floor covering in bedroom1 nd 2. how much floor covering in squre metre will you need?
length of room 1 is 5.0m and width is 5.0
lenth of room 2 is 4.0m and width is 5.0
.
area of room 1 is
5*5 = 25 square metre
area of room 2 is
4*5 = 20 square metre
.
total area = 25+20 = 45 square metre
.
Note: normally, when buying floor covering, you need to ADD additional material to account for "waste" -- typically about 10-15% of the total.